OpenStudy (anonymous):

write the polynomial In completely factored form: f(x)=x^4-16

OpenStudy (gaberen):

By the difference of squares: \[x^4-16=(x^2-4)(x^2+4)=(x-2)(x+2)(x^2+4)\]